# Consumption Partial Insurance

### v3

@@ C.5 Additional results
Notes: The table presents the estimates of the parameters of the linear consumption function, allowing them to vary across wealth and education groups. This table accompanies the quadratic specification estimates in table 6 in the paper. For reasons of brevity, we do not report results when second- and third-order moments of income and consumption are targeted in the estimation; these are available from the authors upon request. Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019. No college versus some college is defined on the basis of the highest level of education attained by the male spouse. All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
@@ Appendix D Welfare costs details
Table D.1 presents the parameter estimates and the model fit of the quantitative model for income shocks (mixture of normal distributions). Table D.2 reports average welfare costs for different values of the CRRA coefficient $\gamma$. Table D.3 calculates the welfare costs respecting the age-specific nature of the partial insurance parameters.
@@
All households 100% -0.72 -0.77 0.66
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.55 6.15 7.48
$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.05 8.36 9.66
$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.16 11.57 12.83
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=1$ (panel A) and $\gamma=3$ (panel B). We simulate consumption for 50,000 households using the time-invariant transmission parameters of table 4. The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
Table D.3: Welfare costs of non-Gaussian relative to Gaussian income risk, age-specific transmission parameters
${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
% households# (1) (2) (3)
All households 100% -1.13 -1.15 -0.24
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.27 5.66 6.51
$\min\{\zeta_{it}\}<-0.70$ 29.11% 5.72 7.71 8.54
$\min\{\zeta_{it}\}<-0.95$ 17.07% 7.85 10.78 11.59
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$. We simulate consumption for 50,000 households using the age-specific transmission parameters of table 5. The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.

### v4

@@ C.6 Additional results
Notes: The table presents the estimates of the parameters of the linear consumption function, allowing them to vary by wealth and education. The underlying income moments also vary by wealth and education (results not shown for brevity). We do not report results when up to third-order moments of income and consumption are targeted in the estimation; these are available upon request. Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019. No college versus some college is defined on the basis of the highest level of education attained by the male spouse (cutoff is 12 years of schooling). All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses. This table accompanies the quadratic specification estimates in table 6 in the paper.
@@ Appendix D Welfare costs details
Table D.1 presents the parameter estimates and model fit of the quantitative model for income (mixture of normal distributions). Table D.2 calculates the welfare costs allowing for heterogeneity in income risk and consumption transmission by age, while table D.3 allows for heterogeneity by wealth. Table D.4 reports average welfare costs for different values of the utility coefficient $\gamma$.
@@
Table D.2: Welfare costs of non-Gaussian relative to Gaussian income risk, age-specific income risk and transmission parameters
${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
% households# (1) (2) (3)
All households 100% -1.02 -1.06 0.31
$\min\{\zeta_{it}\}<-0.50$ 39.18% 3.87 6.07 7.35
$\min\{\zeta_{it}\}<-0.70$ 29.23% 5.14 8.29 9.54
$\min\{\zeta_{it}\}<-0.95$ 18.28% 6.79 11.55 12.75
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$. We simulate consumption for 50,000 households using age-specific income parameters (table C.9) and consumption transmission parameters (table 4). The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
Table D.3: Welfare costs of non-Gaussian relative to Gaussian income risk, by wealth
Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$
${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
% households# (1) (2) (3)
Panel A. Low wealth
All households 100% -1.58 -1.80 6.14
$\min\{\zeta_{it}\}<-0.50$ 40.96% 6.78 7.43 14.64
$\min\{\zeta_{it}\}<-0.70$ 26.87% 9.61 10.63 17.60
$\min\{\zeta_{it}\}<-0.95$ 12.99% 13.12 14.72 21.36
Panel B. High wealth
All households 100% -0.62 -0.69 -0.16
$\min\{\zeta_{it}\}<-0.50$ 47.18% 3.06 4.59 5.09
$\min\{\zeta_{it}\}<-0.70$ 35.12% 4.18 6.42 6.91
$\min\{\zeta_{it}\}<-0.95$ 21.91% 5.73 9.06 9.54
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$ and low (panel A) versus high wealth (panel B). Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 1999–2019. We simulate consumption for 50,000 households using wealth-specific income and consumption transmission parameters, which, for reasons of brevity, are not reported in the main text. However, the consumption parameters can be inferred from table 6, which reports wealth-and-education-specific results, i.e., an even finer cut of the sample. The first row presents the average costs across all households in the wealth group. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
@@
All households 100% -0.81 -0.88 0.66
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.80 6.39 7.82
$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.36 8.65 10.04
$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.54 11.89 13.23
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=1$ (panel A) and $\gamma=3$ (panel B). We simulate consumption for 50,000 households using the time-invariant consumption transmission parameters of table 3. The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.